OpenStudy (anonymous):

Taylor had $126 to spend on 6 books after buying them she has $12 how much did each book cost

Mehek (mehek14):

well here's our equation 6x + 12 = 126

Mehek (mehek14):

first we do is subtract 126 - 12 = ?

OpenStudy (mysticfury):

126 - 12 = 114 divided by 6 = 19 so each book cost 19$

OpenStudy (simplymarie_x):

^

Mehek (mehek14):

@MysticFury it's better to help guide the person instead of just solving and giving them the answer